OpenStudy (anonymous):

Bubble wrap sells in sheets, and each sheet has 1 million individual spherical bubbles, each with a diameter of 0.01 m. What is the total volume of air contained in the bubbles of one sheet of bubble wrap? Round to the nearest tenth.

OpenStudy (anonymous):

@563blackghost

OpenStudy (john_es):

The volume of a spher, \[V=\frac{4\pi r^3}{3}\] Calculate it and then multipli by 1000000.

563blackghost (563blackghost):

So we would basically find the volume of one of the bubbles and then we would multiply it by 1 million.... Formula of a sphere: \(\Huge{V=\frac{4}{3} \pi r^{3}} \times 10^{6} \) So we would first find the radius of the sphere so.... \(\Huge{0.01 \div 2=radius}\)

OpenStudy (anonymous):

Is it .5?

563blackghost (563blackghost):

Do you mean the total volume or the radius?

OpenStudy (anonymous):

volume

563blackghost (563blackghost):

That would be correct ^^

OpenStudy (anonymous):

THanks so much!

563blackghost (563blackghost):

np ^^